9th Standard Syllabus & Materials
9th Standard
TN 9th Tamil உள்ளத்தின் சீர் -கற்கண்டு (இலக்கணம்) - தொடர் இலக்கணம், ஆகுபெயர் Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil உள்ளத்தின் சீர் -விரிவானம் (துணைப்பாடம்) - தாய்மைக்கு வறட்சி இல்லை! Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil உள்ளத்தின் சீர் -கவிதை பேழை (செய்யுள்) - மார்கழி பெருவிழா Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil உயிருக்கு வேர்-இலக்கணம் - பகுபத உறுப்பிலக்கணம் Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil அமுதென்று பேர் - இலக்கணம்-எழுத்து -அளபெடை Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil அமுதென்று பேர்-துணைப்பாடம் - ஆறாம்திணை Important Questions And Answers Study Material - QB365 Set A

Published on: 24/09/2019
Term 3 Probability
Download Tamil Nadu 9th Standard Maths question papers, model tests, one-mark questions, important questions, and public exam papers in PDF format. Free study materials and answer keys for TN State Board students.
Questions + Answers key
Take MCQ Maths Test1.
An urn contains 10 red and 8 white balls. One ball is drawn at random. Find the probability, that the ball drawn is white.
2.
Two dice are thrown simultaneously. Find the probability of getting
(i) an even number as the sum.
(ii) a total of at least 10
(iii) a doublet of even number.
3.
A bag contains 3 red and 2 blue marbles. A marble is drawn at random. What is the probability of drawing a blue marble?
4.
Tickets numbered from 1 to 20 are mixed up together and then a ticket is drawn at random. What is the probability that the ticket has a number which is a multiple of 3 or 7?
5.
What is the probability that a number selected from the numbers 1, 2, 3, ..., 25 is prime number when each of the given numbers is equally likely to be selected?
6.
If a probability of a player winning a particular tennis match is 0.72. What is the probability of the player losing the match?
7.
A company manufactures 10,000 Laptops in 6 months. In that 25 of them are found to be defective. When you choose one Laptop from the manufactured, what is the probability that selected Laptop is a good one.
8.
Frame two problems in calculating probability, based on the spinner shown here.

9.
There are 24 balls in a pot. If 3 of them are Red, 5 of them are Blue and the remaining are Green then, what is the probability of picking out
(i) a Blue ball
(ii) a Red ball and
(iii) a Green ball?
10.
What is the probability of throwing an even number with a single standard dice of six faces?
1.

2.
S = { (1,1), (1,2), (1,3), (1,4), (1,5), (1,6), (2,1), (2,2), (2,3), (2,4), (2,5), (2,6), (3,1), (3,2), (3,3), (3,4), (3,5), (3,6), (4,1), (4,2), (4,3), (4,4), (4,5), (4,6), (5,1), (5,2), (5,3), (5,4), (5,5), (5,6), (6,1), (6,2), (6,3), (6,4), (6,5), (6,6)}
n(S) = 36
(i) Probability of even number as the sum = \(\frac { 18 }{ 36 } =\frac { 1 }{ 2 } \)

3.
Probability = \(\frac{2}{5}\).
4.
A = {3, 6, 9, 12, 15, 18, 7, 14}
P(A) = \(\frac { 8 }{ 20 } =\frac { 2 }{ 5 } \).
5.
A = {2, 3, 5, 7, 11, 13, 17, 19, 23}
P(A) = \(\frac{9}{25}\).
6.
p(A) = 0.72
P(A') = 1 - 0.72 = 0.28
7.
Total n(S) = 10,000
Defective n(A) = 25
Number of defective laptop is 25
Number of good laptop is = 10000 - 25 = 9975
Let A be the event of choosing a good laptop. n(A) = 9975
\(\mathrm{P}(\mathrm{A})=\frac{\mathrm{n}(\mathrm{A})}{\mathrm{n}(\mathrm{S})}=\frac{9975}{10000}\)
P(A) = 0.9975
Probability of getting a good laptop is 0.9975
8.
(i) What is the probability that the spinner will not land on a multiple of 2?
(ii) What is the probability that the spinner will land on an odd number?
9.
n(S) = 24
Red = n(R) = 3
Blue = n(B) = 5
Green= n(G) = 16
(i) Probability of picking a Blue ball =\(\frac { n(B) }{ n(S) } =\frac { 5 }{ 24 } \)
(ii) Probability of picking a Red ball =\(\frac { n(R) }{ n(S) } =\frac { 3 }{ 24 } =\frac { 1 }{ 8 } \)
(iii) Probability of picking a Green ball =\(\frac { n(G) }{ n(S) } =\frac { 16 }{ 24 } =\frac { 2 }{ 3 } \).
10.
Faces of a dice (S) = {1, 2, 3, 4,5, 6}
n(S) = 6
Event of throwing an even number
A = {2, 4, 6}, n(A) = 3
∴ Probability of throwing an even number
P(A) = \(\frac { n(A) }{ n(S) } =\frac { 3 }{ 6 } =\frac { 1 }{ 2 } \).
9th Standard Syllabus & Materials
9th Standard
TN 9th Tamil உள்ளத்தின் சீர் - திருக்குறள் Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil உள்ளத்தின் சீர் - மணிமேகலை Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil உள்ளத்தின் சீர் - ஏறு தழுவுதல் Important Questions And Answers Study Material - QB365 Set A
NEW9th Standard
TN 9th Tamil உயிருக்கு வேர் - தண்ணீர் Important Questions And Answers Study Material - QB365 Set A
Tamilnadu Stateboard 9th Standard Subjects
Tamilnadu Stateboard Standards